### v1

@@ 4.3 Implementation details
All benchmark problems and considered optimizers are implemented on top of the library PyTorch [59]. The code used for generating the numerical results is part of the open-source library DistTraiNN [42]. The description of DNN architectures is summarized in Table 1 and Table 2 for PINNs and DONs, respectively. All networks are initialized using Xavier initialization strategy [22]. For PINNs, the boundary conditions are imposed by employing the length factor function $\ell$ given as $\ell(x_{1},x_{2})=x_{1}(1-x_{1})x_{2}(1-x_{2})$ and $\ell(t,x)=t(x+1)(x-1)$ for DA problem and Burg/AC problems, respectively. Moreover, we employ an adaptive activation function [33] and a skip-connection [28].
@@
As a baseline, we employ the LBFGS optimizer and single-level layer-wise preconditioned LBFGS (SL-BFGS). The proposed two-level preconditioner is also tested by preconditioning the LBFGS method, giving rise to the TL-BFGS optimizer. All LBFGS variants utilize the limited-memory implementation with the memory size $m=3$. Moreover, we utilize momentum with parameter $0.9$ and the cubic backtracking line-search with strong Wolfe conditions [14, Algorithm A6.3.1, pages 325-327] to obtain all step-sizes. For the SL/TL-LBFGS optimizers, the global and local optimization problems outlined in Algorithm 1, are also solved using the LBFGS, which is restarted at every epoch.
@@
Compared to the standard LBFGS algorithm, a single iteration/epoch of the TL-LBFGS method is computationally more expensive. To analyze the iteration cost, Table 3 summarizes the gradient evaluations cost $(g_{e})$, the update cost (UC) and memory cost (MC) of the LBFGS and TL-LBFGS optimizers. As we can see, the LBFGS optimizer requires one gradient evaluation per iteration, i.e., $g_{e}=1$. However, UC depends on the size of stored secant pairs ($m$) used for Hessian approximation. It can be approximately estimated as $4mn$ flops when the compact matrix representation is employed; see [8, Sections 3 and 4] for details. By incorporating the cost associated with evaluating the momentum, the UC equals to $n+4mn$ flops. In the perspective of MC, the LBFGS optimizer requires storing two matrices of size $m\times n$ as well as a momentum of length $n$, which takes up memory of size $n+2mn$.
@@
The computational cost of the TL-LBFGS optimizer is associated with global, local and coarse-level training phases. The global computational cost is exactly same as for the standard LBFGS optimizer. The gradient evaluation cost for local and coarse-level parts are $k_{s}g_{e_{s}}$ and $k_{0}g_{e_{0}}$, where $(g_{e_{s}})$ and $(g_{e_{0}})$ denote the cost of gradient evaluation of local and coarse-level optimizers, respectively. Since the number of coarse trainable parameters is the number of subdomain $N_{sd}$, the cost of coarse-level evaluation $g_{e_{0}}$ is generally less than that of local evaluation $g_{e_{s}}$. In terms of UC, let $\text{UC}_{s}$ and $\text{UC}_{0}$ be the update cost required by the local and coarse optimizers. Since the UC for $s^{th}$ subnetwork is scaled proportionally to the number of local trainable parameters $n_{s}$ and the number of local iterations $k_{s}$, the local update cost is $(\nicefrac{{\tilde{n}_{s}}}{{n}})k_{s}\text{UC}_{s}$. Similarly, the coarse-level update cost is $(\nicefrac{{n_{0}}}{{n}})k_{0}\text{UC}_{0}$. Note that MC of TL-LBFGS can be derived similarly to the derivation of UC, i.e., the local memory cost and the coarse-level memory cost are $(\nicefrac{{\tilde{n}_{s}}}{{n}})\text{MC}_{s}$ and $(\nicefrac{{n_{0}}}{{n}})\text{MC}_{0}$, respectively.
@@
$k_{s}(g_{\text{e}_{s}})+$ $(\nicefrac{{\tilde{n}_{s}}}{{n}})k_{s}\text{UC}_{s}+$ $(\nicefrac{{\tilde{n}_{s}}}{{n}})\text{{MC}}_{s}+$
@@ 5 Numerical results
In this section, we investigate the performance of the proposed TL-LBFGS preconditioner. To this aim, we monitor the value of the loss function $\pazocal{L}$ and the relative $L^{2}$ error, given as
@@ 5.1 Impact of the overlap on the convergence of SL-LBFGS
First, we investigate the impact of overlap on the convergence of the SL-LBFGS optimizer. During this experiment, we fix the number of local iterations at $k_{s}=50$ and consider a varying number of subnetworks $N_{sd}\in\{2,4,8\}$ and the size of the overlapping regions $\delta\in\{0,1,2\}$. Fig. 5 illustrates the obtained results for Burg (PINN) and the Adv (DON) equations. Note, the layer-wise synchronization strategy proposed in Section 3.1.1 is not utilized. Instead, we average the parameters of the layers corresponding to the overlap, an approach traditionally used in numerical analysis. As we can see from the obtained results, in this particular case, increasing the size of the overlap $\delta$ does not improve convergence, and in some cases, it even causes its deterioration.
@@
Figure 6: The convergence of SL-LBFGS method for Burg (Top) and Adv (Bottom) with respect to various $N_{sd}$ and $\delta$. The layer-wise synchronization (LWS) strategy introduced in Section 3.1.1 is utilized. The results are obtained over 5 independent runs.
For comparison, Fig. 6 demonstrates the obtained results by incorporating the novel layer-wise synchronization strategy. In contrast to the results reported in Fig. 5, we can observe that using the layer-wise synchronization strategy allows for an enhanced convergence as the size of the overlapping region $\delta$ increases. Here, we, however, point out that as $\delta$ increases, the computational cost of solving the local minimization problems also increases. Consequently, selecting $\delta=1$ provides the best trade-off between enhanced convergence and the low computational cost.
@@
As a next step, we investigate the impact of coarse-level training on the convergence of the proposed preconditioner. In this experiment, we fix the size of the overlap to $\delta=1$ and the number of local iterations to $k_{s}=50$. At the same time, we vary the number of coarse-level iterations $k_{0}$ and the number of subdomains $N_{sd}$, such that $k_{0}\in\{0,25,50\}$ and $N_{sd}\in\{2,4,8\}$. Figure 7 illustrates the obtained results. As we can see, the loss function decays more rapidly as the number of coarse iterations $k_{0}$ increases, suggesting that incorporating coarse-level training significantly improves the convergence rate of the proposed TL-LBFGS method. However, because the coarse-level training process is sequential and cannot be parallelized, an increased number of coarse-level iterations increases the computational cost. As a consequence, to achieve the tradeoff between the computational cost and the improvement in the convergence speed, we keep $k_{0}$ fairly low for all subsequent numerical experiments, i.e., between $25$ and $50$.
@@
5.3 Convergence, accuracy and computational cost of the LBFGS, and SL/TL-LBFGS optimizers
@@
In this section, we compare the performance of the proposed TL-LBFGS optimizer with the standard LBFGS optimizer. The comparison is performed with respect to relative error $\pazocal{E}_{\text{rel}}$, the number of epochs, $g_{e}$ and UC, as discussed in Section 4.3. Figure 8 and Fig. 9 demonstrate the obtained results for PINN’s and DON’s benchmark problems. As we can see, TL-LBFGS always provides the models with lower $\pazocal{E}_{\text{rel}}$ than the LBFGS optimizer. For example, for the DA example with $N_{sd}=5$, the relative error $\pazocal{E}_{\text{rel}}$ almost immediately decreases to $10^{-2}$, while LBFGS optimizer fails to train the model to that accuracy. Moreover, we can also observe that for all benchmark problems, the performance of TL-LBFGS improves as the number of subdomains $N_{sd}$ increases.
@@
Figure 8: The convergence of LBFGS ($N_{sd}=1$) and TL-LBFGS ($N_{sd}=\{2,5,6\}$) optimizers for training PINN’s benchmark problems with respect to epoch (left), $g_{e}$ (middle), and UC (right). Note that a total of 32, 10, and 8 layers are used for the AC, DA, Burg problems, respectively. For TL-LBFGS, the size of the overlapping region is $\delta=1$. The results are obtained over 5 independent runs.
@@
Figure 9: The convergence of LBFGS ($N_{sd}=1$) and TL-LBFGS ($N_{sd}=\{2,5,6\}$) optimizers for training DON’s benchmark problems with respect to epoch (left), $g_{e}$ (middle), and UC (right). Standard DeepONet is used for Aniso and Helm, while POD-DeepONet is used for Adv. In addition, a total of 10, 12, and 8 layers are used for the Aniso, Helm, and Adv problems, respectively. For TL-LBFGS, the size of the overlapping region is $\delta=1$. The results are obtained over 5 independent runs.
@@
As a next step, we compare the performance of TL-LBFGS with respect to its single-level counterpart, SL-LBFGS. We configure SL-LBFGS as proposed in [40], i.e., without overlap, without layer-wise synchronization strategy and without the coarse-level training. As we can see from Table 4, except in the case of Aniso problem, the TL-LBFGS method consistently outperforms the SL-LBFGS method by providing more accurate DNN models. While the observed improvements compared to the LBFGS method are around one order of magnitude, they are not as big as those compared to the SL-LBFGS method. However, in some cases, such as the Burg benchmark problem, where SL-LBFGS did not perform very well 222 For Burg problem, the SL-LBFGS method starts stagnating after $6{,}000$ epochs at $\pazocal{E}_{\text{rel}}=2\times 10^{-4}$. Therefore, its overall performance to achieve the same $\pazocal{E}_{\text{rel}}$ as the LBFGS method is poor, despite being very efficient in the initial training phase., TL-LBFGS yields about four times more accurate results.
@@
Table 4: Left: The best relative error $\pazocal{E}_{\text{rel}}$ which is achieved by LBFGS, and SL/TL-LBFGS optimizers for PINN’s (AC, DA, Burg) and DON’s (Aniso, Adv, Helm) benchmark problem. The results are obtained over 5 independent runs. Right: An overview of hyperparameters chosen for testing the SL/TL-LBFGS methods. The symbols $N_{sd}$, $k_{s}$, and $k_{0}$ stand for the number of subnetworks, the number of iterations for the local problems, and the number of iterations for the coarse problems, respectively.
@@
In the end, we compare $g_{e}$ and UC required to achieve the same accuracy as the LBFGS optimizer. Tables 5 and 6 report the obtained results in terms of speedup factor $\pazocal{S}$, defined as $\pazocal{S}^{A}_{B}=\nicefrac{{C_{A}}}{{C_{B}}}$, where $C_{A}$ and $C_{B}$ denote the $g_{e}$ or UC required by the optimizer A and by the optimizer B, respectively. As we can see, for PINN’s problems, which are significantly more ill-conditioned than DON’s problems, utilizing the TL-LBFGS optimizer reduces $g_{e}$ and UC by factors of $1.5-1,099$ and $3.5-12,121$ compared to the LBFGS method, respectively. Moreover, compared to SL-LBFGS, we achieve a speedup of $2.9-7.3$ and $1.6-5.1$ for $g_{e}$ and UC, respectively. This effectively highlights the benefits of incorporating the proposed coarse-level training step, overlap, and layer-wise synchronization into the training process.
For DON’s benchmark problems, we observe lower speedup factors than for the PINNs. In particular, compared to LBFGS, we obtain speedup factors of $12.8-18.3$ for $g_{e}$ and $33.2-45.5$ for UC, respectively. Compared to SL-LBFGS, we notice speedup factors of $0.64-1.7$ for $g_{e}$ and $0.3-1.3$ for UC, respectively. The most unsatisfactory result is obtained for the Aniso example, where we even observe an increase in the computational cost. We attribute this behavior to the fact that this problem is the least ill-conditioned of all our benchmark problems, see Appendix A. As a consequence, constructing a two-level preconditioner significantly increases the computational cost, while the convergence speed improves only slightly.
Table 5: The number of gradient evaluation ($g_{e}$) required by LBFGS and SL/TL-LBFGS optimizers for achieving the same relative error $\pazocal{E}_{\text{rel}}$ as obtained by the LBFGS optimizer. The results are obtained over 5 independent runs. The speedup factor $\pazocal{S}$ is defined as $\pazocal{S}^{A}_{B}=\nicefrac{{g_{e}^{A}}}{{g_{e}^{B}}}$, where $g_{e}^{A}$ and $g_{e}^{B}$ denote the $g_{e}$ required by the optimizer A and by the optimizer B.
@@
AC $3.6\times 10^{-3}$ $1{,}090{,}134$ $49{,}296$ $22.11$ $\boldsymbol{16{,}786}$ $64.9$ $2.9$
DA $5.5\times 10^{-1}$ $497{,}000$ $1{,}560$ $318.5$ $\boldsymbol{452}$ $1{,}099.5$ $3.5$
Burg $1.5\times 10^{-4}$ $99{,}900$ $553{,}332$ 0.2 $\boldsymbol{75{,}014}$ 1.5 7.3
Aniso $1.7\times 10^{-2}$ $130{,}623$ $\boldsymbol{4{,}784}$ $27.3$ $7{,}392$ $17.7$ $0.64$
Adv $3.0\times 10^{-2}$ $36{,}338$ $5{,}044$ $7.2$ $\boldsymbol{2{,}838}$ $12.8$ $1.7$
Helm $7.3\times 10^{-2}$ $29{,}965$ $2{,}673$ $11.2$ $\boldsymbol{1{,}638}$ $18.3$ $1.6$
Table 6: The update cost (UC) required by LBFGS and SL/TL-LBFGS optimizers for achieving the same relative error $\pazocal{E}_{\text{rel}}$ as obtained by the LBFGS optimizer. The results are obtained over 5 independent runs. The speedup factor $\pazocal{S}$ is defined as $\pazocal{S}^{A}_{B}=\nicefrac{{\text{UC}^{A}}}{{\text{UC}^{B}}}$, where $\text{UC}^{A}$ and $\text{UC}^{B}$ denote the UC required by the optimizer A and by the optimizer B.

### v2

@@ 4.3 Implementation details
All benchmark problems and considered optimizers are implemented on top of the library PyTorch [65]. The code used for generating the numerical results is part of the open-source library DistTraiNN [66]. The description of DNN architectures is summarized in Table˜1 and Table˜2 for PINNs and DONs, respectively. All networks are initialized using Xavier initialization strategy [67]. For PINNs, the boundary conditions are imposed by employing the length factor function $\ell$ given as $\ell(x_{1},x_{2})=x_{1}(1-x_{1})x_{2}(1-x_{2})$ and $\ell(t,x)=t(x+1)(x-1)$ for DA problem and Burg/AC problems, respectively. Moreover, we employ an adaptive activation function [68] and a skip-connection [43] to prevent gradient vanishing during training and improve the inference performance of the neural network.
@@
As a baseline, we employ the LBFGS optimizer and single-level layer-wise preconditioned LBFGS (SL-LBFGS). The proposed two-level preconditioner is also tested by preconditioning the LBFGS method, giving rise to the TL-LBFGS optimizer. All LBFGS variants utilize the limited-memory implementation with the memory size $m=3$. Moreover, we utilize momentum with parameter $0.9$ and the cubic backtracking line-search with strong Wolfe conditions [69, Algorithm A6.3.1, pages 325-327] to obtain all step-sizes. For the SL/TL-LBFGS optimizers, the global and local optimization problems outlined in Algorithm˜1, are also solved using the LBFGS, which is restarted at every epoch.
@@
Compared to the standard LBFGS algorithm, a single iteration/epoch of the TL-LBFGS method is computationally more expensive. To analyze the iteration cost, Table˜3 summarizes the gradient evaluations cost $(g_{e})$, the update cost (UC) and memory cost (MC) of the LBFGS and TL-LBFGS optimizers. As we can see, the LBFGS optimizer requires one gradient evaluation per iteration, i.e., $g_{e}=1$. However, UC depends on the size of stored secant pairs ($m$) used for Hessian approximation. It can be approximately estimated as $4mn$ flops when the compact matrix representation is employed; see [70, Sections 3 and 4] for details. By incorporating the cost associated with evaluating the momentum, the UC equals to $n+4mn$ flops. In the perspective of MC, the LBFGS optimizer requires storing two matrices of size $m\times n$ as well as a momentum of length $n$, which takes up memory of size $n+2mn$.
@@
The computational cost of the TL-LBFGS optimizer is associated with global, local and coarse-level training phases. The global computational cost is exactly same as for the standard LBFGS optimizer. The gradient evaluation cost for local and coarse-level parts are $k_{s}g_{e_{s}}$ and $k_{0}g_{e_{0}}$, where $(g_{e_{s}})$ and $(g_{e_{0}})$ denote the cost of gradient evaluation of local and coarse-level optimizers, respectively. Since the number of coarse trainable parameters is the number of subdomain $N_{sd}$, the cost of coarse-level evaluation $g_{e_{0}}$ is generally less than that of local evaluation $g_{e_{s}}$. In terms of UC, let $\text{UC}_{s}$ and $\text{UC}_{0}$ be the update cost required by the local and coarse optimizers. Since the UC for $s^{th}$ subnetwork is scaled proportionally to the number of local trainable parameters $n_{s}$ and the number of local iterations $k_{s}$, the local update cost is $(\nicefrac{{\tilde{n}_{s}}}{{n}})k_{s}\text{UC}_{s}$. Similarly, the coarse-level update cost is $(\nicefrac{{n_{0}}}{{n}})k_{0}\text{UC}_{0}$. Similarly, the coarse-level update cost is $(\nicefrac{{n_{0}}}{{n}})k_{0}\,\text{UC}_{0}$. Note that the MC of TL-LBFGS can be derived in the same way as the UC, i.e., the local memory cost and the coarse-level memory cost are $(\nicefrac{{\tilde{n}_{s}}}{{n}})\,\text{MC}_{s}$ and $(\nicefrac{{n_{0}}}{{n}})\,\text{MC}_{0}$, respectively. Moreover, note that in the case of the single-level layer-wise preconditioner for L-BFGS, termed SL-LBFGS, the computational cost can be estimated by simply setting $k_{0}=0$.
Remark 3.
Although the network is decomposed into $N_{sd}$ subnetworks, the application of $\pazocal{M}_{SL/TL}$ requires a full forward/backward pass through the network. This is reflected in the estimate for the number of gradient evaluations ($g_{e}$), which is multiplied by the number of local and coarse-level steps, i.e., by $k_{s}$ and $k_{0}$, respectively. However, the search directions $\{\boldsymbol{d}_{s}\}_{s=1}^{N_{sd}}$ can be computed in parallel, requiring storage only of local gradients and Hessian approximations, as reflected in the estimates for UC and MC.
@@
$k_{s}(g_{\text{e}_{s}})$ $(\nicefrac{{\tilde{n}_{s}}}{{n}})k_{s}\text{UC}_{s}+$ $(\nicefrac{{\tilde{n}_{s}}}{{n}})\text{{MC}}_{s}+$
@@ 5 Numerical results
In this section, we investigate the performance of the proposed layer-wise preconditioner. Since it has been demonstrated in [30, 71] that, for training SciML models, L-BFGS is clearly superior to Adam and other first-order methods, we focus our numerical results on the performance of the two-level layer-wise preconditioner for L-BFGS, giving rise to TL-LBFGS training algorithm. To this aim, we monitor the value of the loss function $\pazocal{L}$ and the relative $L^{2}$ error, given as
@@ 5.1 Impact of the overlap on the convergence of SL-LBFGS
First, we investigate the impact of overlap on the convergence of the SL-LBFGS optimizer. During this experiment, we fix the number of local iterations at $k_{s}=50$ and consider a varying number of subnetworks $N_{sd}\in\{2,4,8\}$ and the size of the overlapping regions $\delta\in\{0,1,2\}$. Figure˜6 illustrates the obtained results for Burg (PINN) and the Adv (DON) equations. Note, the layer-wise synchronization strategy proposed in Section˜3.1.1 is not utilized. Instead, we average the parameters of the layers corresponding to the overlap, an approach traditionally used in numerical analysis. As we can see from the obtained results, in this particular case, increasing the size of the overlap $\delta$ does not improve convergence, and in some cases, it even causes its deterioration.
@@
Figure 7: The convergence of SL-LBFGS method for Burg (Top) and Adv (Bottom) with respect to various $N_{sd}$ and $\delta$. The layer-wise synchronization (LWS) strategy introduced in Section˜3.1.1 is utilized. The results are obtained over 5 independent runs.
For comparison, Figure˜7 demonstrates the obtained results by incorporating the novel layer-wise synchronization strategy. In contrast to the results reported in Figure˜6, we can observe that using the layer-wise synchronization strategy allows for an enhanced convergence as the size of the overlapping region $\delta$ increases. Here, we, however, point out that as $\delta$ increases, the computational cost of solving the local minimization problems also increases. Consequently, selecting $\delta=1$ provides the best trade-off between enhanced convergence and the low computational cost.
@@
As a next step, we investigate the impact of coarse-level training on the convergence of the proposed preconditioner. In this experiment, we fix the size of the overlap to $\delta=1$ and the number of local iterations to $k_{s}=50$. At the same time, we vary the number of coarse-level iterations $k_{0}$ and the number of subdomains $N_{sd}$, such that $k_{0}\in\{0,25,50\}$ and $N_{sd}\in\{2,4,8\}$. Figure˜8 illustrates the obtained results. As we can see, the loss function decays more rapidly for a sufficient number of coarse iterations $k_{0}$, suggesting that incorporating coarse-level training significantly improves the convergence rate of the proposed TL-LBFGS method. However, because the coarse-level training process is sequential and cannot be parallelized, an increased number of coarse-level iterations increases the computational cost. Moreover, we would like to remark that the coarse-level model is guaranteed to approximate the fine-level model only in the local neighborhood of the current fine-level iterate and is therefore inherently local. Consequently, performing too many $k_{0}$ nonlinear steps on the coarse level may cause the coarse-level model to no longer be consistent with the fine-level model, which can in turn degrade overall convergence. To balance computational cost with improvements in convergence speed, we keep $k_{0}$ fairly low for all subsequent numerical experiments, i.e., between $25$ and $50$.
@@
5.3 Convergence, accuracy and computational cost of LBFGS and SL/TL-LBFGS optimizers
In this section, we compare the performance of the proposed TL-LBFGS optimizer with the standard LBFGS optimizer. The comparison is performed with respect to relative error $\pazocal{E}_{\text{rel}}$, the number of epochs, $g_{e}$ and UC, as discussed in Section˜4.3. Our first experiment illustrates the impact of preconditioning on the overall performance of training. We compare the performance of the two-level additive Schwarz method as a standalone solver (TL-SA) and as a preconditioner for LBFGS (TL-LBFGS). The obtained results clearly indicate that the preconditioned variant is significantly more efficient. These observations are consistent with findings reported in the literature. For example, V. Dolean et al. [38] and X. Cai et al. [37] demonstrated that the nonlinear two-level additive Schwarz method alone is insufficient to achieve fast convergence. To overcome this limitation, they used the two-level additive Schwarz to precondition the Newton method. In view of the superiority of the preconditioning strategy, all remaining experiments employ the proposed two-level additive Schwarz method as a preconditioner for L-BFGS.
Figure 9: The convergence of two-level additive Schwarz method (TL-SA) used as an standalone solver and as a preconditioner for LBFGS (TL-LBFGS). The symbol $\pazocal{L}(\boldsymbol{\theta})$ denotes the training loss.
@@ 5.3 Convergence, accuracy and computational cost of LBFGS and SL/TL-LBFGS optimizers
Figure˜10 and Figure˜11 compare the performance of the LBFGS and TL-LBFGS for PINN’s and DON’s benchmark problems. As we can see, TL-LBFGS always provides the models with lower $\pazocal{E}_{\text{rel}}$ than the LBFGS optimizer. For example, for the DA example with $N_{sd}=5$, the relative error $\pazocal{E}_{\text{rel}}$ almost immediately decreases to $10^{-2}$, while LBFGS optimizer fails to train the model to that accuracy. Moreover, we can also observe that for all benchmark problems, the performance of TL-LBFGS improves as the number of subdomains $N_{sd}$ increases.
@@
Figure 10: The convergence of LBFGS ($N_{sd}=1$) and TL-LBFGS ($N_{sd}=\{2,4\}$} for AC, $N_{sd}=\{2,5\}$ for DA, and $N_{sd}=\{2,6\}$ for Burg) optimizers for training PINN’s benchmark problems with respect to epoch (left), $g_{e}$ (middle), and UC (right). Note that a total of 32, 10, and 8 layers are used for the AC, DA, Burg problems, respectively. For TL-LBFGS, the size of the overlapping region is $\delta=1$. The results are obtained over 5 independent runs.
@@
Figure 11: The convergence of LBFGS ($N_{sd}=1$) and TL-LBFGS ($N_{sd}=\{2,4\}$} for Adv, $N_{sd}=\{2,5\}$ for Aniso, and $N_{sd}=\{2,6\}$ for Helm) optimizers for training DON’s benchmark problems with respect to epoch (left), $g_{e}$ (middle), and UC (right). Standard DeepONet is used for Aniso and Helm, while POD-DeepONet is used for Adv. In addition, a total of 10, 12, and 8 layers are used for the Aniso, Helm, and Adv problems, respectively. For TL-LBFGS, the size of the overlapping region is $\delta=1$. The results are obtained over 5 independent runs.
@@
As a next step, we compare the performance of TL-LBFGS with respect to its single-level counterpart, SL-LBFGS. We configure SL-LBFGS as proposed in [30], i.e., without overlap, without layer-wise synchronization strategy and without the coarse-level training. As we can see from Table˜4, except in the case of Aniso problem, the TL-LBFGS method consistently outperforms the SL-LBFGS method by providing more accurate DNN models. While the observed improvements compared to the LBFGS method are around one order of magnitude, they are not as big as those compared to the SL-LBFGS method. However, in some cases, such as the Burg benchmark problem, where SL-LBFGS did not perform very well  [^3] For Burg problem, the SL-LBFGS method starts stagnating after 6,0006{,}000 epochs at $\pazocal{E}_{\text{rel}}=2\times 10^{-4}$. Therefore, its overall performance to achieve the same $\pazocal{E}_{\text{rel}}$ as the LBFGS method is poor, despite being very efficient in the initial training phase., TL-LBFGS yields about four times more accurate results.
@@
Table 4: Left: The best relative error $\pazocal{E}_{\text{rel}}$ which is achieved by LBFGS, and SL/TL-LBFGS optimizers for PINN’s (Allen-Cahn equation (AC), Diffusion-Advection (DA), and Burgers’ (Burg)) and DON’s (Anisotropic Poisson (Aniso), Advection (Adv), and Helmholtz (Helm)) benchmark problem. The results are obtained over 5 independent runs. Right: An overview of hyperparameters chosen for testing the SL/TL-LBFGS methods. The symbols $N_{sd}$, $k_{s}$, and $k_{0}$ stand for the number of subnetworks, the number of iterations for the local problems, and the number of iterations for the coarse problems, respectively. Note that $N_{sd}$ takes two different values, reflecting the use of two distinct decompositions. For example, $2,3$ indicate that $N_{sd}=2$ and 33 was used.
@@
In the end, we compare $g_{e}$ and UC required to achieve the same accuracy as the LBFGS optimizer. Tables 5 and 6 report the obtained results in terms of speedup factor $\pazocal{S}$, defined as $\pazocal{S}^{A}_{B}=\nicefrac{{C_{A}}}{{C_{B}}}$, where $C_{A}$ and $C_{B}$ denote the $g_{e}$ or UC required by the optimizer A and by the optimizer B, respectively. As we can see, for PINN’s problems, which are significantly more ill-conditioned than DON’s problems, utilizing the TL-LBFGS optimizer reduces $g_{e}$ and UC by factors of 1.51,0991.5-1,099 and 3.512,1213.5-12,121 compared to the LBFGS method, respectively. Moreover, compared to SL-LBFGS, we achieve a speedup of 2.97.32.9-7.3 and 1.65.11.6-5.1 for $g_{e}$ and UC, respectively. This effectively highlights the benefits of incorporating the proposed coarse-level training step, overlap, and layer-wise synchronization into the training process.
For DON’s benchmark problems, we observe lower speedup factors than for the PINNs. In particular, compared to LBFGS, we obtain speedup factors of 12.818.312.8-18.3 for $g_{e}$ and 33.245.533.2-45.5 for UC, respectively. Compared to SL-LBFGS, we notice speedup factors of 0.641.70.64-1.7 for $g_{e}$ and 0.31.30.3-1.3 for UC, respectively. The most unsatisfactory result is obtained for the Aniso example, where we even observe an increase in the computational cost. We attribute this behavior to the fact that this problem is the least ill-conditioned of all our benchmark problems, see A. As a consequence, constructing a two-level preconditioner significantly increases the computational cost, while the convergence speed improves only slightly.
Table 5: The number of gradient evaluation ($g_{e}$) required by LBFGS and SL/TL-LBFGS optimizers for achieving the same relative error $\pazocal{E}_{\text{rel}}$ as obtained by the LBFGS optimizer for PINN’s (Allen-Cahn (AC), Diffusion-Advection (DA), and Burgers’ (Burg)) and DON’s (Anisotropic Poisson (Aniso), Advection (Adv), and Helmholtz (Helm)) benchmark problem. The results are obtained over 5 independent runs. The speedup factor $\pazocal{S}$ is defined as $\pazocal{S}^{A}_{B}=\nicefrac{{g_{e}^{A}}}{{g_{e}^{B}}}$, where $g_{e}^{A}$ and $g_{e}^{B}$ denote the $g_{e}$ required by the optimizer A and by the optimizer B.
@@
AC $3.6\times 10^{-3}$ 1,090,1341{,}090{,}134 49,29649{,}296 $22.11$ $\boldsymbol{16{,}786}$ $64.9$ $2.9$
DA $5.5\times 10^{-1}$ 497,000497{,}000 1,5601{,}560 $318.5$ $\boldsymbol{452}$ 1,099.51{,}099.5 $3.5$
Burg $1.5\times 10^{-4}$ 99,90099{,}900 553,332553{,}332 0.2 $\boldsymbol{75{,}014}$ 1.5 7.3
Aniso $1.7\times 10^{-2}$ 130,623130{,}623 $\boldsymbol{4{,}784}$ $27.3$ 7,3927{,}392 $17.7$ $0.64$
Adv $3.0\times 10^{-2}$ 36,33836{,}338 5,0445{,}044 $7.2$ $\boldsymbol{2{,}838}$ $12.8$ $1.7$
Helm $7.3\times 10^{-2}$ 29,96529{,}965 2,6732{,}673 $11.2$ $\boldsymbol{1{,}638}$ $18.3$ $1.6$
Table 6: The update cost (UC) required by LBFGS and SL/TL-LBFGS optimizers for achieving the same relative error $\pazocal{E}_{\text{rel}}$ as obtained by the LBFGS optimizer for PINN’s (Allen-Cahn (AC), Diffusion-Advection (DA), and Burgers’ (Burg)) and DON’s (Anisotropic Poisson (Aniso), Advection (Adv), and Helmholtz (Helm)) benchmark problem. The results are obtained over 5 independent runs. The speedup factor $\pazocal{S}$ is defined as $\pazocal{S}^{A}_{B}=\nicefrac{{\text{UC}^{A}}}{{\text{UC}^{B}}}$, where $\text{UC}^{A}$ and $\text{UC}^{B}$ denote the UC required by the optimizer A and by the optimizer B.